# Enhancing LLMs for Governance with Human Oversight: Evaluating and Aligning LLMs on Expert Classification of Climate Misinformation for Detecting False or Misleading Claims about Climate Change

### v1

@@ Abstract
Climate misinformation is a problem that has the potential to be substantially aggravated by the development of Large Language Models (LLMs). In this study we evaluate the potential for LLMs to be part of the solution for mitigating online dis/misinformation rather than the problem. Employing a public expert annotated dataset and a curated sample of social media content we evaluate the performance of proprietary vs. open source LLMs on climate misinformation classification task, comparing them to existing climate-focused computer-assisted tools and expert assessments. Results show (1) state-of-the-art (SOTA) open-source models substantially under-perform in classifying climate misinformation compared to proprietary models, (2) existing climate-focused computer-assisted tools leveraging expert-annotated datasets continues to outperform many of proprietary models, including GPT-4o, and (3) demonstrate the efficacy and generalizability of fine-tuning GPT-3.5-turbo on expert annotated dataset in classifying claims about climate change at the equivalency of climate change experts with over 20 years of experience in climate communication. These findings highlight 1) the importance of incorporating human-oversight, such as incorporating expert-annotated datasets in training LLMs, for governance tasks that require subject-matter expertise like classifying climate misinformation, and 2) the potential for LLMs in facilitating civil society organizations to engage in various governance tasks such as classifying false or misleading claims in domains beyond climate change such as politics and health science.
@@ 1 Introduction
When considering public support for climate change science and policy, people rely on information from mediated sources such as online or offline media rather than directly from scientists Scheufele (2014). The reliance on such sources provide an opportunity for false or misleading claims about climate change to compete with accurate ones and influence public opinion and policy discourse which seriously hampers climate mitigation efforts Allgaier (2019); Gounaridis and Newell (2024); IPCC (2022); Lewandowsky (2021); Treen et al. (2020).
@@
Advances in Large Language Models (LLMs) could also contribute to the exacerbation and diffusion of climate misinformation if exploited by malicious actors. LLMs are capable of generating high volumes of persuasive, deceptive, and human-like content full of misinformation that promotes climate change denial and skepticism Fore et al. (2024); Zhang et al. (2024); CAAD (2024); Ellison and Hugh (2024) making it more difficult for humans to detect LLM-generated dis/misinformation in news context especially when circulating on social media Kreps et al. (2022); Marlow et al. (2020).
@@
The major knowledge gap about the nature of climate change information on digital platforms, combined with inconsistent and ineffective content moderation policies across platforms CAAD (2024); CCDH (2021); Romero-Vicente (2023), continue to seriously hamper the effective mitigation of climate change misinformation on social media. A key element of this problem is the inadequate identification and classification tools that capture the technical expertise required to evaluate the veracity of claims about climate change circulating online Coan et al. (2021); Vu et al. (2023).
@@
In response, researchers have attempted to develop a variety of LLM-assisted classification tools to identify and respond to false or misleading claims about climate change Coan et al. (2021); Leippold et al. (2024). However, the generalizability of these detectors is less robust on types and sources of text beyond what they are trained on, demonstrating how easy it is for malicious actors to overcome the detection by these algorithms Stiff and Johansson (2022). Additionally, existing LLM benchmarks do not always represent real-word tasks Ni et al. (2024) and typically frame climate misinformation detection as a binary problem Thulke et al. (2024); Lacombe et al. (2023). Existing evaluations of LLMs are predominantly centered around detecting the presence of climate misinformation in text Thulke et al. (2024), rather than on identifying the types of claims that are used to convey such misinformation. Thus, these benchmarks fall short from guiding domain and policy experts in selecting the most suitable model for detecting and classifying claims about climate change at a scale.
@@
To address these issues, we first benchmark the performance of 11 open-source and 5 proprietary (i.e., closed-source) models, on a zero-shot task, in classifying false or misleading claims (FMC) about climate change using a public dataset on climate-misinformation Coan et al. (2021). This dataset, sourced from contrarian and skeptical domains about climate change, also includes the annotation of these claims by climate-communication experts, which serve as the ground truth for comparing LLM responses. Next, we selected the best performing LLM and took a human-centered approach into evaluating the generalizability of the LLM in detecting false or misleading claims about climate change present in articles sourced from low credible news sources that are circulating on social media. We incorporate the assessment of two experts with over 20 years of experience in climate communication into the evaluation of the capability of these LLMs in classifying claims that form the basis of climate misinformation. By pairing a quantitative evaluation with expert assessment of LLMs, we overcome some of the limitations related to the validity, quality, and diversity of automated benchmarks that are commonly used by the AI and NLP communities when performing automated evaluation of LLMsGehrmann et al. (2023); Xiao et al. (2024).
@@ 3. RQ3:
    Does fine-tuning on expert-annotate dataset enhances the model’s ability to generalize better and classify false or misleading claims about climate change circulating on social media?
@@ 4. RQ4:
The results of our study provide the following four contributions: (1) Although open-source LLMs perform well on evaluation benchmarks, we found that they substantially underperform compared to proprietary models in accurately classifying false or misleading claims about climate change on an expert-annotated dataset. (2) We observe and report an inferior performance of proprietary LLMs (GPT-4o, GPT-4o-mini, GPT-4, and Gemini-1.5-flash) as compared to a BERT-based computer assisted tool (CARDS) Coan et al. (2021) for classifying false or misleading claims about climate change. (3) We demonstrate the superior performance of fine-tuning GPT-3.5-turbo, as compared to other proprietary models and CARDS Coan et al. (2021), for classifying false or misleading claims commonly found in climate skeptic and contrarian blogs as annotated by climate change experts. (4) We illustrate GPT-3.5-turbo’s strong and extensible capability for classifying false or misleading claims about climate change commonly found in social media from sources beyond conservative think tanks and contrarian blogs with the approximate reliability as two senior climate change communication experts with over 20 years of experience. Through our findings, we also report a skewed distribution of the types of claims classified by the LLMs potentially reflecting an underlying bias in their training data.
@@ 2 Related Work
The potential abuse of LLMs by malign actors to generate misinformation that shapes the information environment and public opinion has become evident across a range of domains from politics to healthcare, and including climate change Yang and Menczer (2024); Marlow et al. (2020); Ferrara et al. (2020); Akhtar et al. (2023); De Angelis et al. (2023). However, detecting climate dis/misinformation generated by humans is difficult, let alone content generated by LLMs Chen and Shu (2023). In an effort to combat climate dis/misinformation, researchers have introduced tools that leverage LLMs to identify false claims about climate change from skeptic and contrarian sources, fact-check these claims, and even generate factual and in-depth answers to climate related questions Coan et al. (2021); Leippold et al. (2024); Thulke et al. (2024); Mullappilly et al. (2023). Despite these efforts, researchers have questioned the generalizability of dis/misinformation detection tools and have identified vulnerabilities that enable malicious content to bypass them Stiff and Johansson (2022). One proposed solution to this issue is to establish benchmarks to evaluate the capability of LLMs in classifying and detecting climate related content.
Researchers have constructed several datasets relevant to climate change research, but not always tailored for benchmarking the detection of specific types of climate dis/misinformation. Existing datasets are being used to evaluate LLMs for climate-related classification tasks such as: predicting sentence relevance to climate change Leippold et al. (2024), stance detection in support or opposition toward climate change prevention Vaid et al. (2022), and fact-checking scientific information related to climate science Laud et al. (2023); Pirozelli et al. (2023).
Still, such benchmarks seem to be suffering from limitations in terms of the diversity Ni et al. (2024), complexity Liang et al. (2022), and representation of how climate dis/misinformation is being manifested in terms of false or misleading claims Thulke et al. (2024). It is essential, therefore, to involve stakeholders such as climate change experts in the design, development, and validation for many of these tools and datasets, as their input and feedback not only refines the quality of the benchmark datasets, but also contributes to model enhancements in terms of handling climate misinformation Stiennon et al. (2020); Zhou and Xu (2020); Ouyang et al. (2022); Christiano et al. (2017). Through such human-centered approach, stakeholders can identify in-depth criteria for what these models are being evaluated on and propose edge cases that impose constraints on the model performance before it gets deployed or leveraged for governance tasks Xiao et al. (2024).
@@ 3 Data
Our study employs two datasets: 1) a public dataset of false or misleading claims about climate change used to train Climate Change Denial and Skepticism (CARDS) model Coan et al. (2021) and 2) a curated sample of the most engaging articles and blog posts (i.e., based on the number of likes, comments, shares) about climate change on Facebook and X (i.e., formerly known as Twitter) from right-biased, questionable, and low credible sources.
@@ 3.1 Evaluating LLMs using CARDS
The curated dataset to train the CARDS model contains paragraphs in English language from articles sourced from 53 contrarian and skeptical domains about climate change spanning the years 1998 to 2020 and their corresponding claims based on annotations from the authors who are experts in climate research Coan et al. (2021). The dataset is randomly split into training (N=23,436), validation (N=2,605), and test sets (N=2,904). Our decision to use this dataset for evaluating LLMs is based on the (1) breadth of this dataset on climate misinformation and (2) the taxonomy of claims developed and validated by experts in climate-research Coan et al. (2021). In addition, the availability of the taxonomy of super-claims and coding manuals used to annotate the claims also enables us to craft prompts using the instructions in these manuals as part of the zero-shot task to classify false and misleading claims about climate change (as later described in section 4.1). The CARDS taxonomy of super claims and sub-claims represents the primary arguments employed by climate denialists and skeptics (see Appendix A.2). False and misleading claims in this taxonomy are grouped into five main categories: (1) global warming is not happening, (2) humans greenhouse gases are not causing global warming, (3) climate impacts are not bad, (4) climate solutions won’t work, and (5) the climate movement and/or science are unreliable. Claims in the CARDS dataset are formatted as strings that combine the super-claim and sub-claim into a single label separated by an underscore (e.g., "5_1") referring to the super-claim and sub-claim, respectively.
@@ 3.2 Social media dataset
Domain credibility. We then filtered our corpus by the domain credibility of the publisher as means to curate a test dataset that contained a sufficient number of false or misleading claims about climate change circulating on social media. To determine the domain level credibility of each scraped article, we relied on a combination of the Media Bias Fact Check (MBFC) categories MediaBiasFactCheck (2024) and NewsGuard Trust score NewsGuard (2024). MBFC categorizes news sources in one of nine bias categories: least biased, left bias, left-center bias, right-center bias, right bias, conspiracy-pseudoscience, questionable sources, pro-science, and satire. Similarly, NewsGuard Trust score is a reliability rating between 0 and 100 that is assigned by journalists and editors to news websites based on journalistic and apolitical criteria such as credibility and transparency. A NewsGuard score of 60 or below indicate an untrustworthy news source.
@@ 4.1 Benchmarking LLMs using CARDS
To assess how well open-source vs. proprietary LLMs perform in classifying false or misleading claims about climate change, we apply zero-shot classification technique to classify paragraphs about climate change, sourced from articles published by conservative think tanks and blog posts, that are in the CARDS test dataset. This dataset serves as a baseline for comparing the claims classified by each LLM with the annotated claims by climate-research experts in the CARDS test dataset on the paragraph level. All evaluations of the 11 open-source models in this research were conducted on Google Colab, which constrained our model selection process for this task based on the availability of the compute and memory resources. A full list of the LLMs used for this assessment and their performance can be found in Table 2. Open-source models in the aforementioned list were selected based on the limited computing resources we had access to on Google Colab. In contrast, for closed-source models, we rely on OpenAI models and Gemini because these models are widely used in various applications, including climate research Zhu and Tiwari (2023); Kraus et al. (2023); Mullappilly et al. (2023). This limitation reflects a similar constraint faced by civil organizations interested in this research and application area but lacking the resources to host high-performing open-source models. Along with base models, we chose to included instruction-tuned models because they are generally aligned to mitigate social harms, such as misinformation and manipulation, and tend to demonstrate improved reasoning Achiam et al. (2023); Dubey et al. (2024), which may translate to a better performance on our benchmarking task.
@@
For closed source models, we included a system and user prompts (A.4 and A.5, respectively) based on the instructions derived from the coding manual used to train annotators for labeling the training and testing datasets that were subsequently used to train the RoBERTalarge CARDS model Coan et al. (2021).
Similarly, we structured the prompts for the open-source models to reflect a question-answer prompt template as illustrated in A.8. This structure is commonly used for evaluating LLMs in zero-shot tasks Rajpurkar et al. (2018).
@@
Although we prompted the models at the sub-claim level to match CARDS training dataset’s conceptual framework structured around specific subclaims, as outlined in the original CARDS paper Coan et al. (2021). However, for analysis and applications aimed at refuting climate misinformation, we will focus only at the super-claim level. Accordingly, to assess the performance of the different models in classifying super-claims, we compare the classified claims using the LLMs with the annotated claims in the CARDS test dataset at the super-claim level of analysis. To this end, we split the LLM-generated claim labels and those in the test dataset based on the under score delimiter. Only the number to the left of the underscore ("_") delimiter was extracted for model evaluation and comparison as it represents the super-claim. We elaborate on our findings from benchmarking the performance of the different LLMs with respect to the test dataset of the CARDS model in section 5.1 of the results section.
@@
The need to constantly update, test, and validate prompts, especially for emerging claims, makes zero-shot prompting a less scalable approach when compared to fine-tuning. Fine-tuning enhances the ability of general purpose LLMs to be more optimal in their adaptability to domain-specific tasks such as identifying emerging types of claims about climate change by including a few examples of these claims in the training data. In addition, fine-tuned LLMs that accurately classify false or misleading information about climate change have the potential to be integrated into a variety of governance tools such as those used for content moderation Leippold et al. (2024).
For these reasons, along with our findings from benchmarking different LLMs as elaborated in section 5.1 of results, we chose to fine-tune GPT-3.5-turbo [^3] Fine-tuning GPT-4o was not yet available for fine-tuning at the time of conducting this analysis to classify and compare the classified false or misleading claims about climate change of the fine-tuned model with the claims classified by the other models in the zero-shot task described in section 4.1, and those annotated in the CARDS test set by climate-research experts.
Employing OpenAI’s Python library, we fine-tune GPT-3.5-turbo for 3 epochs with a batch size of 8 using the training (N=23,436) and validation (N=2,605) datasets from the CARDS. The number of epochs and batch size were selected to closely resemble the training hyperparamters for RoBERTalarge CARDS model [^4] RoBERTalargeCARDS model was trained on 3 epochs and batch size of 6.

### v2

@@ Abstract
Climate misinformation is a problem that has the potential to be substantially aggravated by the development of Large Language Models (LLMs). In this study we evaluate the potential for LLMs to be part of the solution for mitigating online dis/misinformation rather than the problem. Employing a public expert annotated dataset and a curated sample of social media content we evaluate the performance of proprietary vs. open source LLMs on climate misinformation classification task, comparing them to existing climate-focused computer-assisted tools and expert assessments. Results show (1) open-source models substantially under-perform in classifying climate misinformation compared to proprietary models, (2) existing climate-focused computer-assisted tools leveraging expert-annotated datasets continue to outperform many of proprietary models, including GPT-4o, and (3) demonstrate the efficacy and generalizability of fine-tuning GPT-3.5-turbo on expert annotated dataset in classifying claims about climate change at the equivalency of climate change experts with over 20 years of experience in climate communication. These findings highlight 1) the importance of incorporating human-oversight, such as incorporating expert-annotated datasets in training LLMs, for governance tasks that require subject-matter expertise like classifying climate misinformation, and 2) the potential for LLMs in facilitating civil society organizations to engage in various governance tasks such as classifying false or misleading claims in domains beyond climate change such as politics and health science.
keywords:
[…]
@@ 1 Introduction
When considering public support for climate change science and policy, people rely on information from mediated sources such as online or offline media rather than directly from scientists [1]. The reliance on such sources provide an opportunity for false or misleading claims about climate change to compete with accurate ones and influence public opinion and policy discourse which seriously hampers climate mitigation efforts [2, 3, 4, 5, 6].
@@
Advances in Large Language Models (LLMs) could also contribute to the exacerbation and diffusion of climate misinformation if exploited by malicious actors. LLMs are capable of generating high volumes of persuasive, deceptive, and human-like content full of misinformation that promotes climate change denial and skepticism [7, 8, 9, 10] making it more difficult for humans to detect LLM-generated dis/misinformation in news context especially when circulating on social media [11, 12].
@@
The major knowledge gap about the nature of climate change information on digital platforms, combined with inconsistent and ineffective content moderation policies across platforms [9, 13, 14], continue to seriously hamper the effective mitigation of climate change misinformation on social media. A key element of this problem is the inadequate identification and classification tools that capture the technical expertise required to evaluate the veracity of claims about climate change circulating online [15, 16].
@@
In response, researchers have attempted to develop a variety of LLM-assisted classification tools to identify and respond to false or misleading claims about climate change [15, 17]. However, the generalizability of these detectors is less robust on types and sources of text beyond what they are trained on, demonstrating how easy it is for malicious actors to overcome the detection by these algorithms [18]. Additionally, existing LLM benchmarks do not always represent real-word tasks [19] and typically frame climate misinformation detection as a binary problem [20, 21]. Existing evaluations of LLMs are predominantly centered around detecting the presence of climate misinformation in text [20], rather than on identifying the types of claims that are used to convey such misinformation. Thus, these benchmarks fall short from guiding domain and policy experts in selecting the most suitable model for detecting and classifying claims about climate change at a scale.
@@
To address these issues, we first benchmark the performance of 11 open-source and 5 proprietary (i.e., closed-source) models, on a zero-shot task, in classifying false or misleading claims (FMC) about climate change using a public dataset on climate-misinformation [15]. This dataset, sourced from contrarian and skeptical domains about climate change, also includes the annotation of these claims by climate-communication experts, which serve as the ground truth for comparing LLM responses. Next, we selected the best performing LLM and took a human-centered approach into evaluating the generalizability of the LLM in detecting false or misleading claims about climate change present in articles sourced from low credible news sources that are circulating on social media. We incorporate the assessment of two experts with over 20 years of experience in climate communication into the evaluation of the capability of these LLMs in classifying claims that form the basis of climate misinformation. By pairing a quantitative evaluation with expert assessment of LLMs, we overcome some of the limitations related to the validity, quality, and diversity of automated benchmarks that are commonly used by the AI community when performing automated evaluation of LLMs [22, 23].
@@ 3. RQ3:
    Does fine-tuning on expert-annotated dataset enhance the model’s ability to generalize better and classify false or misleading claims about climate change circulating on social media?
@@ 4. RQ4:
The results of our study provide the following four contributions: (1) Although open-source LLMs perform well on evaluation benchmarks, we found that they substantially underperform compared to proprietary models in accurately classifying false or misleading claims about climate change on an expert-annotated dataset. (2) We observe and report an inferior performance of proprietary LLMs (GPT-4o, GPT-4o-mini, GPT-4, and Gemini-1.5-flash) as compared to a BERT-based computer assisted tool (CARDS) [15] for classifying false or misleading claims about climate change. (3) We demonstrate the superior performance of fine-tuning GPT-3.5-turbo, as compared to other proprietary models and CARDS [15], for classifying false or misleading claims commonly found in climate skeptic and contrarian blogs as annotated by climate change experts. (4) We illustrate GPT-3.5-turbo’s strong and extensible capability for classifying false or misleading claims about climate change commonly found in social media from sources beyond conservative think tanks and contrarian blogs with the approximate reliability as two senior climate change communication experts with over 20 years of experience. Through our findings, we also report a skewed distribution of the types of claims classified by the LLMs potentially reflecting an underlying bias in their training data.
@@ 2 Related Work
The potential abuse of LLMs by malign actors to generate misinformation that shapes the information environment and public opinion has become evident across a range of domains from politics to healthcare, and including climate change [24, 12, 25, 26, 27]. However, detecting climate dis/misinformation generated by humans is difficult, let alone content generated by LLMs [28]. In an effort to combat climate dis/misinformation, researchers have introduced tools that leverage LLMs to identify false claims about climate change from skeptic and contrarian sources, fact-check these claims, and even generate factual and in-depth answers to climate related questions [15, 17, 20, 29]. Despite these efforts, researchers have questioned the generalizability of dis/misinformation detection tools and have identified vulnerabilities that enable malicious content to bypass them [18]. One proposed solution to this issue is to establish benchmarks to evaluate the capability of LLMs in classifying and detecting climate related content.
Researchers have constructed several datasets relevant to climate change research, but not always tailored for benchmarking the detection of specific types of climate dis/misinformation. Existing datasets are being used to evaluate LLMs for climate-related classification tasks such as: predicting sentence relevance to climate change [17], stance detection in support or opposition toward climate change prevention [30], and fact-checking scientific information related to climate science [31, 32].
Still, such benchmarks seem to be suffering from limitations in terms of the diversity [19], complexity [33], and representation of how climate dis/misinformation is being manifested in terms of false or misleading claims [20]. It is essential, therefore, to involve stakeholders such as climate change experts in the design, development, and validation for many of these tools and datasets, as their input and feedback not only refines the quality of the benchmark datasets, but also contributes to model enhancements in terms of handling climate misinformation [34, 35, 36, 37]. Through such human-centered approach, stakeholders can identify in-depth criteria for what these models are being evaluated on and propose edge cases that impose constraints on the model performance before it gets deployed or leveraged for governance tasks [23].
@@ 3 Data
Our study employs two datasets: 1) a public dataset of false or misleading claims about climate change used to train Climate Change Denial and Skepticism (CARDS) model [15] and 2) a curated sample of the most engaging articles and blog posts (i.e., based on the number of likes, comments, shares) about climate change on Facebook and X (i.e., formerly known as Twitter) from right-biased, questionable, and low credible sources.
@@ 3.1 Evaluating LLMs using CARDS
The curated dataset to train the CARDS model contains paragraphs in English language from articles sourced from 53 contrarian and skeptical domains about climate change spanning the years 1998 to 2020 and their corresponding claims based on annotations from the authors who are experts in climate research [15]. The dataset is randomly split into training (N=23,436), validation (N=2,605), and test sets (N=2,904). Our decision to use this dataset for evaluating LLMs is based on the (1) breadth of this dataset on climate misinformation and (2) the taxonomy of claims developed and validated by experts in climate-research [15]. In addition, the availability of the taxonomy of super-claims and coding manuals used to annotate the claims also enables us to craft prompts using the instructions in these manuals as part of the zero-shot task to classify false and misleading claims about climate change (as later described in section 4.1). The CARDS taxonomy of super claims and sub-claims represents the primary arguments employed by climate denialists and skeptics (see Appendix A.2). False and misleading claims in this taxonomy are grouped into five main categories: (1) global warming is not happening, (2) humans greenhouse gases are not causing global warming, (3) climate impacts are not bad, (4) climate solutions won’t work, and (5) the climate movement and/or science are unreliable. Claims in the CARDS dataset are formatted as strings that combine the super-claim and sub-claim into a single label separated by an underscore (e.g., "5_1") referring to the super-claim and sub-claim, respectively.
@@ 3.2 Social media dataset
Domain credibility. We then filtered our corpus by the domain credibility of the publisher as means to curate a test dataset that contained a sufficient number of false or misleading claims about climate change circulating on social media. To determine the domain level credibility of each scraped article, we relied on a combination of the Media Bias Fact Check (MBFC) categories [38] and NewsGuard Trust score [39]. MBFC categorizes news sources in one of nine bias categories: least biased, left bias, left-center bias, right-center bias, right bias, conspiracy-pseudoscience, questionable sources, pro-science, and satire. Similarly, NewsGuard Trust score is a reliability rating between 0 and 100 that is assigned by journalists and editors to news websites based on journalistic and apolitical criteria such as credibility and transparency. A NewsGuard score of 60 or below indicate an untrustworthy news source.
@@ 4.1 Benchmarking LLMs using CARDS
To assess how well open-source vs. proprietary LLMs perform in classifying false or misleading claims about climate change, we apply zero-shot classification technique to classify paragraphs about climate change, sourced from articles published by conservative think tanks and blog posts, that are in the CARDS test dataset. This dataset serves as a baseline for comparing the claims classified by each LLM with the annotated claims by climate-research experts in the CARDS test dataset on the paragraph level. All evaluations of the 11 open-source models in this research were conducted on Google Colab, which constrained our model selection process for this task based on the availability of the compute and memory resources. A full list of the LLMs used for this assessment and their performance can be found in Table 1. Open-source models in the aforementioned list were selected based on the limited computing resources we had access to on Google Colab. In contrast, for closed-source models, we rely on OpenAI models and Gemini because these models are widely used in various applications, including climate research [40, 41, 29]. This limitation reflects a similar constraint faced by civil organizations interested in this research and application area but lacking the resources to host high-performing open-source models. Along with base models, we chose to included instruction-tuned models because they are generally aligned to mitigate social harms, such as misinformation and manipulation, and tend to demonstrate improved reasoning [42, 43], which may translate to a better performance on our benchmarking task.
@@
For closed source models, we included a system and user prompts (A.4 and A.5, respectively) based on the instructions derived from the coding manual used to train annotators for labeling the training and testing datasets that were subsequently used to train the RoBERTalarge CARDS model [15].
Similarly, we structured the prompts for the open-source models to reflect a question-answer prompt template as illustrated in A.8. This structure is commonly used for evaluating LLMs in zero-shot tasks [44].
@@
Although we prompted the models at the sub-claim level to match CARDS training dataset’s conceptual framework structured around specific subclaims, as outlined in the original CARDS paper [15]. However, for analysis and applications aimed at refuting climate misinformation, we will focus only at the super-claim level. Accordingly, to assess the performance of the different models in classifying super-claims, we compare the classified claims using the LLMs with the annotated claims in the CARDS test dataset at the super-claim level of analysis. To this end, we split the LLM-generated claim labels and those in the test dataset based on the under score delimiter. Only the number to the left of the underscore ("_") delimiter was extracted for model evaluation and comparison as it represents the super-claim. We elaborate on our findings from benchmarking the performance of the different LLMs with respect to the test dataset of the CARDS model in section 5.1 of the results section.
@@
The need to constantly update, test, and validate prompts, especially for emerging claims, makes zero-shot prompting a less scalable approach when compared to fine-tuning. Fine-tuning enhances the ability of general purpose LLMs to be more optimal in their adaptability to domain-specific tasks such as identifying emerging types of claims about climate change by including a few examples of these claims in the training data. In addition, fine-tuned LLMs that accurately classify false or misleading information about climate change have the potential to be integrated into a variety of governance tools such as those used for content moderation [17].
For these reasons, along with our findings from benchmarking different LLMs as elaborated in section 5.1 of results, we chose to fine-tune GPT-3.5-turbo [^3] Fine-tuning GPT-4o was not yet available for fine-tuning at the time of conducting this analysis. to classify and compare the classified false or misleading claims about climate change of the fine-tuned model with the claims classified by the other models in the zero-shot task described in section 4.1, and those annotated in the CARDS test set by climate-research experts.
Employing OpenAI’s Python library, we fine-tune GPT-3.5-turbo for 3 epochs with a batch size of 8 using the training (N=23,436) and validation (N=2,605) datasets from the CARDS. The number of epochs and batch size were selected to closely resemble the training hyperparamters for RoBERTalarge CARDS model [^4] RoBERTalargeCARDS model was trained on 3 epochs and batch size of 6..